07 March 2008

RHIC Running:  Cryo test yesterday we lowered the flow rate of liquid helium to observe the 10 Hz orbit oscillations.  Results are not conclusive.  It appears that the oscillation frequency is changed, perhaps to random, but the amplitude of the oscillation is about the same.  The cryo test was performed by reducing the flow rate to 100 g/s, which gives a temperature increase between recoolers of 100mK.  The flow rate can be dropped to 60 g/s with a delta-T of 190mK.  The 10Hz study could be repeated at the 60 g/s flow rate to look for more definitive results.

There will be an attempt to use the JET to measure both Blue and Yellow simultaneously. 

Fill 9992 had polarization of 56/53 at injection, dropping to 56/45 at store.  At lower intensity (1.5/1.4e11) the Yellow ramp efficiency has recovered to 98%. 

Carbon Polarimeter measurement seems to fail frequently.  There needs to be a new set of instructions, and improved reliability in the Controls Software.

Mei Bai is working with Todd to perform a polarization measurement on the Ramp. 

Waldo will continue to do snake scans.

PHENIX is interested in two vernier scans; one long at the end of a store. 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020